how do you do this problem? i just need an example and im on my way!
40% of 50=
thank you

Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

40% of 50=
-----------
per cent means "out of 100", 1 cent = 1/100 of a dollar, century = 100 years, etc
-----------
40% = 40/100 = 0.4
0.4*50 = 20

Answer by richwmiller(17219)

" of " means multiply.
40%=.40=4/10=2/5 4/10 or 2/5 are the easiest forms to use here
.40*50=20.00=20
2/5*50=2*10=20
4/10*50=4*5=20